Here is some old data I showed Linda's group on 4/17 started at 16h02 am on
6/15/04 and lasting about 42.6 hrs until 6/17 at 10h39am. There were no top up
failures during this time. At the end of this time serie, we found out that the
chopper horizontal position had moved and we turned off the motor. These time
series for plotted to make sure the beam did not move on the slit too much.
Attachment 5 shows the intensities. Experiments are ungoing so access to 7ID-C
sends the BPM values to large negative values. Linda's group scanned the
monochromator energy several times, detuned the mono 2nd crystal several time serie.
Attachment 1 shows the 7ID-C X-ray BPM data. Note that the vertical beam
moves by about 150 um peak to peak, so one expect the transmitted intensity to
vary quite a bit.
Attachment 2 shows the APS XBPM data projected to the L5-20 plane. Nothing much
diffrent than the last few days. The APS BPM vertical motion is less noticable
with the feedback on.
Attachment 3 shows the cryocooler information.
Attachment 4 shows the piezo voltage, showing when a several tweaks happened
during several energy scans of the mono.
Tonight at 17h35 (6/18) I started a new time series. Sorry I was too distracted
yesterday and I did not restart the time series.